# Measuring AI Ability to Complete Long Tasks

### v2

@@ 8 Discussion
8.1 Measuring and interpreting time horizon
Context and skill effects
At real companies, junior software engineer hires often take weeks of onboarding to begin contributing economic value. The human baseliners that determine the length of our tasks have much less context than average employees, potentially increasing measured task length. Our tasks are designed to require minimal context, which somewhat mitigates this problem; our internal PRs (Section 6.4) were not designed this way, and so baseliners took many times longer than employees. However, highly skilled baseliners can also complete tasks far faster than average employees. Our expert baseliners are likely much more skilled than the average software engineer, potentially decreasing our measured task length.
Task distribution effects
Measuring extreme time horizons
AGI will have “infinite” horizon length
An infinite time horizon does not mean an arbitrarily capable AI, merely the ability to complete tasks that take humans an arbitrarily long length of time. If an artificial general intelligence (AGI) is capable of completing all tasks expert humans can with a success rate of at least X%, its Xtime horizon will necessarily be infinite. Therefore, if such systems are ever developed, the long-term trend in time horizon will be faster than exponential, with an asymptote at the date of AGI deployment.
Human time horizon measurements
In theory, one could also measure the time horizon of a human or population of humans. However, there are both theoretical and practical difficulties to doing so. We discuss this more in Appendix B.1.1.
8.2 Limitations and future work
More models with better elicitation
In general, we find that properly eliciting models can make a very large difference in their performance. [^19] This is a common obervation; see e.g. the improvements to software development capabilities from AIDE [42] or METR’s recent work on KernelBench: https://metr.org/blog/2025-02-14-measuring-automated-kernel-engineering/. We have put a limited amount of effort into eliciting models to get good performance on our tasks, so while our results are a reasonable lower bound, some models may have somewhat greater capabilities than we demonstrate. The most work was done to elicit o1 and the original Claude 3.5 Sonnet, each of which had around 2-3 engineer weeks of iterative development. All other models use the same scaffolding with at most minor changes. Future work could replicate our results with more effort spent on eliciting the full capabilities of frontier models.
More rigorous human baselining
Our per-task human time estimates are likely noisy due to relatively small sample size, and potentially also systematically skewed in various ways. Most notably, we select only successful completions of a task, and encourage baseliners to give up on tasks they may not complete in a reasonable amount of time. Our baseliners’ skills also vary significantly, and a wide variety of skills are relevant to our tasks. Though we attempt to match baseliners with appropriate tasks, this process is unlikely to be perfect. From manual reviews of baseline attempts, we also observe that humans sometimes simply give up even when the task seems within their capabilities, and it is unclear what selection effects are produced on the distribution of success times as a result. See See Appendix B.1.1 for further discussion of how bias in human baseline times could affect our results. Future work could replicate our results with more rigorous human baseliner selection or explore how sensitive the results are to methodological choices around human baselining.
More natural, varied tasks
There are reasons to believe that our task distribution is systematically different from the distribution of economically valuable work (and perhaps systematically different than the distribution of risk-model relevant tasks). We explored some of these reasons in Sections 6, but there remain many differences that we did not explore. For example, the modality of interaction in our tasks is also relatively narrow—for example, none of our tasks require the use of a mouse. None of our tasks require cooperating or competing with humans or other agents, [^20] See Xu et al. [43] for an example of a recent benchmark that requires multi-agent interaction in a relatively realistic setting. while real software engineering or ML research involves communicating and coordinating with managers and other engineers or researchers. Many real-world tasks require very high reliability, and these are underrepresented in our dataset due to the difficulty of measuring models on these tasks. Most importantly, the tasks we study are heavily skewed toward software engineering and ML research. Future work could explore how the capabilities of AI agents are progressing in other domains.
More use of inference compute
Our scaffolds made relatively limited use of inference-time compute. When assuming that the human expert is compensated at \$143.61/hour (average L4 Engineer salary at Google divided by 2,000 hours), more than 80% of successful runs cost less than 10% of what it would cost for a human to perform the same task. (Figure 13). This implies that if inference-time computation could be used to improve performance, there is substantial room to do so while still remaining economically competitive with human experts. Previous research has found that techniques such as best-of-k can substantially improve performance on a subset of these tasks,[2] and better use of inference-time compute may lead to substantially different scores.
Data analysis
8.3 Summary
@@
In this paper, we proposed an intuitive, quantitative metric for AI capabilities: the task completion time horizon, which relates AI performance on tasks to the typical length of time human experts require to complete the tasks. We constructed a dataset of 66 shorter SWAA tasks, combined these with tasks from RE-Bench and HCAST, and conducted 236 human runs on SWAA tasks to estimate the difficulty of the tasks, combining these difficulty estimates with baselines collected from RE-Bench and HCAST. To measure the trend in time horizon, we benchmarked 11 frontier AI models released between 2019 and 2025 on our dataset, calculated the time horizon of each model (Section 4), and then plotted this against release date.
We observed that the 50% task completion time horizon on our tasks has been growing exponentially from 2019–2025 with a doubling time of approximately seven months (Figure 1), a similar trend to our exploratory work on the trend since 2023 without SWAA data (Section 6.1). Measuring the 80% time horizon revealed a similar exponential trend, though these horizons are approximately 5x shorter than the 50% horizons (Section 4.2.1). Our qualitative analysis (Section 5) identified several factors driving this progress. We also noted important limitations of current systems, particularly their lower performance on less structured, “messier” tasks (Section 6.2).
To investigate the extent to which our observed trend is externally valid (Section 6), we replicated our methods on SWE-bench Verified (Section 6.3) and analyzed the impact of task “messiness”” on model performance (Section  6.2). We observed a similar exponentially increasing time horizon for both SWE-Bench and subsets of our tasks categorized by low and high messiness. However, due to systematic differences between these benchmarks and real-world tasks, these results may still not generalize to actual real-world tasks.
@@
Finally, we attempt to extrapolate the trend on our tasks to one-month (167 hours) AI (Section 7.1), finding that if both the trend continues and observed performance trends generalize to real-world tasks, an 80% confidence interval for the release date of AI that can complete 1-month long software tasks spans from late 2028 to early 2031 (Section 7.2).
Acknowledgments
@@
Appendix A Task suite details
@@
RE-Bench
Figure 15: The 7 original RE-Bench tasks.
HCAST
Details about HCAST tasks are provided in Rein et al. [2025].
A.1 SWAA
@@
Appendix B Methodological details
B.1 Human baselines
For information about HCAST baselines, see Rein et al. [2025].
RE-Bench baselines
@@ B.1 Human baselines
SWAA baselines
Unlike the HCAST and RE-Bench sets, SWAA is baselined by METR employees with relevant expertise using a custom webapp that enables more accurate timing. Because these tasks are intended to be a single step and exclude context acquisition, the timer for SWAA tasks starts after reading general instructions and ends as soon as the user chooses a response, so that baseline time only includes reading the question itself and choosing an answer. Baselines done with our usual setup have timing overhead of seconds to tens of seconds, which would be unacceptably high for single-step tasks with duration shorter than 2 seconds.
Results
We use 286 successful baselines out of 454 attempts on the HCAST and RE-Bench tasks. Across the 149 tasks in this document, 133 have their duration computed from baselines.
B.1.1 Baseline success rate
@@
Figure 16: Success rates and time horizon of human baseliners. Note that the time horizon is not directly comparable to the time horizon of ML models (see Section B.1.1)
B.1.2 Baseline time analysis
@@
B.2 Internal PR tasks details
We collected five issues from an internal repository to better validate model performance on real world tasks.
@@ B.2 Internal PR tasks details
We collected five real, recent, uncontaminated issues from an internal METR repository. We then ran GPT-4o, Claude 3.5 Sonnet (New), and o1 on these issues. We recorded how long it took repository maintainers to solve these issues, and also had external baseliners attempt to fix these issues. Unlike SWE-bench Verified, we did no filtering to require these issues to be automatically verifiable. Instead, maintainers of the relevant repository scored model and baseliner solutions manually, as if they were reviewing PRs:
@@ B.3 How models were run
B.4 Item Response Theory
Item response theory (IRT) is a collection of statistical techniques used to predict the probability that a person will answer a question correctly, based on their ability level. The standard IRT three parameter model is (from Cai et al.):
$T_{i}(1|\eta)=\gamma_{i}+\frac{1-\gamma_{i}}{1+\exp[-(\alpha_{i}+\beta_{i}\eta)]}$
Where $T_{i}(1|\eta)$ indicates the probability that a person with ability level $\eta$ gets question $i$ correctly. IRT is usually used to regress on both the question difficulty and the ability level simultaneously; in contrast, we define the question difficulty as the logarithm of human baseline time, and therefore only regress on ability level. In particular, we set
$T_{i}(1|model)=\sigma((\log h_{model}-\log t_{task})\cdot\beta_{model})$
which has the following changes from the IRT model:
@@ B.4 Item Response Theory
    $\gamma_{i}=0$, indicating that the probability of correctly guessing the answer is zero. (This is effectively true for most tasks, but some SWAA tasks are multiple-choice with a 25% chance of guessing correctly. This choice only affects the time horizon of GPT-2 and GPT-3 and does not meaningfully affect our results.)
@@
    $\alpha_{i}=\log t_{task}$ is calculated with $t_{task}$ as the geometric mean of human baseline times.
@@
    $\eta=\log h_{model}$ is the log 50% time horizon of the model rather than an abstract ability parameter
  4. 4.
    $\beta_{model}$ is an agent-dependent learned parameter, rather than being task-dependent
and regress to determine $\eta$ for each agent. We weight tasks by diversity, such that a task from a family of size $n$ gets weight $1/\sqrt{n}$. We then set a time horizon for each model (model_horizon), which corresponds to the time at which $prob(success)=0.5$. Note that since $logistic(0)=0.5$, this is equivalent to finding the baseline length $\beta_{i}$ such that $\alpha-\beta_{i}\eta=0$, or equivalently $\alpha=\beta_{i}\eta$. For example, following this methodology o1 has a time horizon of about 39 minutes (see Figure 5).
Appendix C Qualitative analysis examples
We provide examples of both ways that models have improved, as well as their major limitations, as described in Section 5
C.1 Ways in which models have improved

### v3

@@ 6 Discussion
Summary
@@
In this paper, we proposed an intuitive, quantitative metric for AI capabilities: the task completion time horizon, which relates AI performance on tasks to the typical length of time human experts require to complete the tasks. We constructed a dataset of 66 shorter SWAA tasks, combined these with existing benchmarks, and conducted baselines to estimate their difficulty. To measure the trend in time horizon, we benchmarked 11 frontier AI models released between 2019 and 2025 on this dataset, and found (Section 3) that the 50% task completion time horizon on these tasks has been growing exponentially from 2019–2025 with a doubling time of approximately seven months (Figure 1). We also noted important limitations of current systems, particularly their lower performance on less structured, “messier” tasks (Appendix F.2).
@@
Finally, we attempt to extrapolate the trend on these tasks to one-month (167 hours) AI (Section 5), finding that if the trend continues and observed performance trends generalize to real-world tasks, an 80% confidence interval for the release date of AI that can complete 1-month long software tasks spans from mid-2028 to mid-2030 (Section 5)– or even as soon as early 2027 if the 2024–2025 trend continues.
Limitations and future work
Even though we believe the exponential trend to be robust, we caution that time horizon is always measured relative to a domain, task distribution, and human baseliners’ level of skill and context, and can be difficult to measure in practice, especially for longer horizons or success rates near 100%. We discuss nuances with interpreting time horizon in Appendix E.3. Finally, in Appendix E.4, we discuss ways of extending or improving our results, such as by studying multiple domains, more models, better elicitation, inference scaling, or other statistical methods.
Acknowledgments and Disclosure of Funding
@@
Appendix A Expanded Related work
Below, we expand on our discussion of related work in Section 1.1.
A.1 Agent and capability benchmarks
The evaluation of AI capabilities has evolved significantly from single-task benchmarks to complex, multi-step evaluations designed to assess agent-like behavior. While traditional benchmarks such as GLUE [Wang et al., 2018], SuperGLUE [Wang et al., 2019], and MMLU [Hendrycks et al., 2020] have provided valuable insights into language model performance, they primarily measure static knowledge rather than the dynamic problem-solving capabilities essential for real-world applications. Recent work has developed more complex agent benchmarks. AgentBench [Liu et al., 2023] evaluates agents across diverse environments including web browsing, coding, and game playing. MLAgentBench [Huang et al., 2024] and MLEBench [Chan et al., 2024] focus specifically on machine learning research tasks, while ToolBench [Qin et al., 2023, Guo et al., 2024] assesses tool use capabilities. The recent ZeroBench [Roberts et al., 2025] involve difficult reasoning, but in the context of visual puzzles rather than economically valuable tasks. Other noteworthy benchmarks include GAIA [Mialon et al., 2024], which evaluates reasoning across multiple modalities, and BIG-bench [Srivastava and others, 2022], which contains hundreds of diverse tasks including many requiring multi-step reasoning.
Software engineering has emerged as a particularly informative domain for evaluating AI capabilities. HumanEval [Chen et al., 2021] and MBPP [Austin et al., 2021] provide programming challenges of varying complexity, while more complex benchmarks like SWE-bench [Jimenez et al., 2024] and APPS [Hendrycks et al., 2021] test more sophisticated programming abilities. We use SWE-bench Verified [Chowdhury et al., 2024]’s human time estimates for task completion in our work. RE-Bench [Wijk et al., 2024], which we use in this work, evaluates models on complex research engineering tasks that may require hours of human effort and compares AI performance to human machine learning engineers. SWE-Lancer [Miserendino et al., 2025] generates tasks
A.2 Forecasting AI progress
Several recent efforts have been made to contextualize AI benchmark performance. One such effort is the annual AI Index Report [Maslej et al., 2024], which tracks performance across various benchmarks, including the date at which models achieved human-level performance. Murray et al. [2025] had cybersecurity experts relate AI performance on Cybench [Zhang et al., 2024] to the ability to autonomously develop malware. Phuong et al. [2024] commissioned professional forecasters to predict whether AI ranks amongst top public concerns by 2030, conditioned on benchmark performance. These efforts generally lack a unified, quantitative metric for cross-benchmark comparison.
A.3 Psychometric methods and Item Response Theory
Our methodological approach draws inspiration from psychometric testing, particularly Item Response Theory (IRT) [Baker, 2001], which models the relationship between latent traits (such as ability) and observed responses to test items. In traditional IRT, item difficulty is a parameter in a logistic model predicting response correctness based on respondent ability. Our approach inverts this, using task completion time (a proxy for difficulty) to predict AI performance. Our methodology also relates to difficulty estimation techniques in educational testing [de Ayala, 2017], where multiple metrics including completion time are used to estimate the difficulty of tasks. IRT has been applied to machine learning classifiers by Martínez-Plumed et al. [2019], and was used to design efficient benchmarks in Song and Flach [2021].
Appendix B Task suite details
As with most benchmarks, the three task suites we study were designed to isolate a specific unit of work that can be reliably accomplished within a time limit. This usually means that the tasks require much less context than the average task in the middle of a larger project. [^10] We define context as information that experienced employees use to complete a task which is not explicitly in the task description or possessed by most external experts. For example, when fixing a bug in a software package, the package’s maintainer may use their experience with past bugs in the same package to guess at the bug’s cause, fluency with the codebase to find the bug, and knowledge of their organization’s priorities to decide whether to apply a quick patch or a more thorough fix. We discuss possible effects of context in Section 4. We confirmed that all tasks were possible given the instructions provided by having humans successfully complete each task at least once. [^11] As many of these attempts were done by in-house staff or were done using different methodology, we exclude these attempts from our human baseline numbers.
B.1 Task subsuites
B.1.1 HCAST suite
Each task is automatically scored between 0 and 1, with higher scores indicating better performance. Many tasks only return scores of 0 or 1, but for tasks with continuous scoring, we manually define a success threshold score, which we use in some of our analysis to binarize agent scores. [^12] A small subset of these tasks is available on request. (We do not share the content of most tasks to reduce the likelihood of AI systems accidentally or intentionally being trained on them.)
More details about HCAST tasks are provided in Rein et al. [2025].
B.1.2 RE-Bench suite
B.1.3 Software atomic actions (SWAA) suite
Example SWAA task (file selection) In contrast to other simple benchmarks like LAMBADA [Paperno et al., 2016] or GSM8K [Cobbe et al., 2021], which test skills not directly applicable to software engineering, the SWAA set represents actions that are needed in both software engineering work and the longer tasks we study. SWAA consists of 5 task families, three representing common decisions, one for code completion, and one for math; see Appendix B.1.3 for more details. Development of SWAA tasks was blind to AI agent performance; that is, all tasks were written before seeing AI attempts, and elicitation (development of the few-shot prompt used for evaluation) was carried out on a separate development task suite.
@@ B.1.3 Software atomic actions (SWAA) suite
RE-Bench
Figure 8: The 7 original RE-Bench tasks.
B.2 Limitations of the task suite
The tasks we use to benchmark AI capabilities are systematically different from real tasks. These differences could result in the trends we observe on these tasks not generalizing to real world tasks. For instance, all SWAA, HCAST, and RE-Bench differ from real-world tasks in the following ways:
  • No interaction with other agents No tasks involve interacting with other autonomous agents. Coordinating with, or competing with, other agents seems likely to increase task difficulty. For instance, by increasing the importance of strategic decision-making, real-time coordination, and predicting the actions of other complex agents.
  • Lax resource constraints No SWAA tasks, and few HCAST tasks saliently involve making efficient use of a limited resource—a common constraint in real-world tasks.
  • Unpunishing Similarly, very few tasks are punishing of single mistakes. [^14] With the exception of submitting an answer too early on HCAST tasks. This is in part to reduce the expected cost of collecting human baselines. Real world tasks can often be more punishing, for instance, when they involve competing against other agents. For instance, a single blunder in a chess game can greatly reduce the chance of winning the game.
  • Static environments Tasks in this suite typically use environments that do not significantly change unless directly acted upon by the agent. In contrast, real tasks often occur in the context of a changing environment.
We attempted to measure how these systematic differences might affect AI agent performance in Section F.2, by including the above properties as “messiness” factors. We found that though the absolute performance on “messier” tasks was lower, the trends in performance were similar to less messy tasks.
Appendix C Methodological details
C.1 Human baselines
C.1.1 HCAST tasks
We use existing baselines collected as part of HCAST. These baselines are collected from domain professionals with relevant experience in software engineering, ML, and cybersecurity. Baseliners work in the same environment as agents, using Vivaria (an open source platform for language model agent evals) with their screens and audio recorded for manual review, to prevent cheating. They are incentivized with bonuses for successful completion and for completing tasks faster than other baseliners. After screening out attempts on tasks other than our subset of HCAST, failed attempts, and those with issues (such as using disallowed AI tools), we include 286 successful baselines from approximately 460 total attempts. Task durations are calculated using the geometric mean of successful baselines, with manual estimates for tasks lacking successful baselines. [^15] Non-baseline estimates were based on information including the length of QA runs that did not follow strict baseline conditions, and the length of similar tasks with successful baselines.
For further information about HCAST baselines, see Rein et al. [2025].
C.1.2 RE-Bench baselines
For RE-Bench, we used the baselines from [Wijk et al., 2024]. As baseliners were instructed to achieve the best performance for each task, we consider the task duration of each of these 6 tasks as 8 hours, and instead use the mean score achieved by baseliners who spent between 7 and 9 hours to convert raw score into success threshold.
@@ C.1.2 RE-Bench baselines
C.1.3 SWAA baselines
Unlike HCAST and RE-Bench, which were baselined by external contractors, SWAA is baselined by internal employees with relevant expertise using a custom webapp that enables more accurate timing. Because these tasks are intended to be a single step and exclude context acquisition, the timer for SWAA tasks ends as soon as the user chooses a response. For decision based tasks, only one selection is allowed to avoid random guessing; for fill in the blank tasks, baseliners can try multiple times until getting the answer correctly or opting to skip. We baselined each decision-based task 4 times and each fill-in-the-blank style task 3 times.
Unlike the HCAST and RE-Bench sets, SWAA is baselined by internal employees with relevant expertise using a custom webapp that enables more accurate timing. Because these tasks are intended to be a single step and exclude context acquisition, the timer for SWAA tasks starts after reading general instructions and ends as soon as the user chooses a response, so that baseline time only includes reading the question itself and choosing an answer. Baselines done with our usual setup have timing overhead of seconds to tens of seconds, which would be unacceptably high for single-step tasks with duration shorter than 2 seconds.
C.1.4 Baseline success rate
@@
Figure 9: Success rates and time horizon of human baseliners. Note that the time horizon is not directly comparable to the time horizon of ML models (see Section C.1.4)
C.1.5 Baseline time analysis
@@
C.2 Internal PR tasks
Summary
We ran GPT-4o, Claude 3.5 Sonnet (New), and o1 on five uncontaminated issues from an internal repository. Resolving these issues was real work performed by internal staff, so we might expect results on these tasks to better represent performance on real economically valuable tasks than a typical benchmark task.
We find that our contract baseliners take 5x-18x longer to resolve issues than repository maintainers. Additionally, AI agent performance on these issues is not inconsistent with AI agent success rate curves derived from HCAST, SWAA, and RE-Bench performance if contractor time-to-complete is used to measure the tasks length. However, it takes much longer for our contract baseliners to complete these tasks than repository maintainers. This suggests that time horizons may have better correspondence to the labor of a low-context human, rather than a high-context human.
@@ C.2 Internal PR tasks
We collected five real, recent, uncontaminated issues from an internal repository. We then ran GPT-4o, Claude 3.5 Sonnet (New), and o1 on these issues. We recorded how long it took repository maintainers to solve these issues, and also had external baseliners attempt to fix these issues. Unlike SWE-bench Verified, we did no filtering to require these issues to be automatically verifiable. Instead, maintainers of the relevant repository scored model and baseliner solutions manually, as if they were reviewing PRs:
@@ C.3 How models were run
All models were run using agent scaffolds using the Anthropic or OpenAI APIs.
C.3.1 Scaffold and platform details
Most AI models were evaluated with modular-public—a basic agent scaffold. This scaffold provides the model with Python and Bash commands and some very simple context management to keep the input within the context window length of the LM. We used a slightly different scaffold for o1-preview and o1, because they seemed to struggle with tool use, responding to environmental feedback, and generally acting as an agent.
@@ C.3.1 Scaffold and platform details
C.3.2 Compute usage
The total amount of compute used included about 2,000 H100-hours for RE-Bench environments and 50,000 CPU hours for other environments on a combination of cloud and internal machines, plus roughly 50,000 H100-hour equivalents of compute used internally or from API providers for inference. This includes compute used for preliminary experiments. Compute used for data analysis is minimal.
C.4 Comparison to Item Response Theory
Item response theory (IRT) is a collection of statistical techniques used to predict the probability that a person will answer a question correctly, based on their ability level. The standard IRT two-parameter logistic (2PL) model is (simplified from Cai et al.):
$T_{i}(1|\eta_{agent})=\sigma((\eta_{agent}-\alpha_{task})\cdot\beta_{task})$
Where $T_{i}(1|\eta_{agent})$ indicates the probability that a person with ability level $\eta_{agent}$ answers a task correctly.  [^16] In the three-parameter logistic model, a parameter $\gamma_{task}$ is added representing the probability of correctly guessing the task. In the 2PL model and our method, $\gamma_{task}=0$. This is effectively true for most tasks, but some SWAA tasks are multiple-choice with a 25% chance of guessing correctly. This choice only affects the time horizon of GPT-2 and GPT-3 and does not meaningfully affect our results. IRT is usually used to regress on both the question difficulty and the ability level simultaneously; in contrast, we define the question difficulty as the logarithm of human baseline time, and therefore only regress on ability level. In particular, our model is as follows.
$T_{i}(1|h_{agent})=\sigma((\log h_{agent}-\log t_{task})\cdot\beta_{agent})$ (1)
Note the following changes from the 2PL model:
@@ C.4 Comparison to Item Response Theory
    $\alpha_{task}=\log t_{task}$ where $t_{task}$ is the geometric mean of successful human baseline times.
@@
    $\eta_{agent}$ can now be interpreted as $\log h_{agent}$ where $h_{agent}$ is the 50% time horizon of the agent rather than an abstract ability parameter. This is because when $h_{agent}=t_{task}$, our model predicts a success probability of $\sigma(0\cdot\beta_{agent})=\sigma(0)=0.5$.
@@
    $\beta_{agent}$ is an agent-dependent learned parameter, rather than being task-dependent.
To fit equation 1, we weight tasks by diversity, such that a task from a family of size $n$ gets weight $1/\sqrt{n}$. We then use logistic regression to fit the time horizon $h_{agent}$ and slope $\beta_{agent}$ for each model. For example, following this methodology o1 has a time horizon of about 39 minutes (see Figure 4).
These changes allow us to compute an interpretable time horizon without several complexities of IRT. First, IRT has an initial calibration/joint optimization stage in which the task difficulty is estimated from all agent scores, so the ability score of one agent may depend on other agents’ performance; in our method each agent’s time horizon is computed individually. Second, IRT outputs ”ability scores” which then need to be converted to a time horizon estimate; our method eliminates that step by estimating horizon directly. Third, the 2PL model uses one slope parameter per task, whereas we use one slope parameter per agent. In our setting with many fewer agents than tasks, this reduces the overall number of parameters.
In exchange, we forego the ability to use IRT metrics designed for test design, such as Fisher information functions for individual tasks (used in e.g. Kipnis et al. [2025]). Future improved methodologies could include these metrics to design sample-efficient benchmarks and evaluations grounded in human data.
Appendix D Qualitative analysis
D.1 Failure categories
We report the results in Table 3.3. We find that over a third of the GPT-4 failures resulted from repeating failed actions, compared to 2 out of 32 for o1, which we see as quantitative evidence for our claim that models seem to have improved in their ability to adapt to mistakes. Half of the o1 failures resulted from abandoning the task prematurely, while only a quarter of the GPT-4 failures resulted from the same—this may result from o1 failures occurring on qualitatively more difficult tasks, or may reflect idiosyncrasies of o1.
D.2 Specific skills of newer models
To better contextualize the observed trend of improved model performance, we examined the transcripts for the tasks where earlier models (e.g. GPT-4) do substantially worse than current models. Specifically, we categorized all task families based on the type of expertise required, and noticed that current models tend to outperform earlier models for tasks that involved ML training, reverse engineering compiled software binaries, and cybersecurity CTFs. In addition, we noticed that agents improved greatly on tasks requiring situational awareness of the AI agent’s limitations or defeating an opposing strategy. This gave us five sets of task families to examine.
For each of these task family sets, we then used contractors to manually read through all runs from all models for all tasks in each of these five task families, and identified possible explanations for the improvement in AI agent performance on those tasks, as well as potential limitations.
We provide examples of both ways that models have improved, as well as their major limitations, as described in Section 3.3.
D.2.1 What do newer models do better?